# Linguistic and Embedding-Based Profiling of Texts generated by Humans and Large Language Models

### v2

@@ Abstract
The rapid advancements in large language models (LLMs) have significantly improved their ability to generate natural language, making texts generated by LLMs increasingly indistinguishable from human-written texts. While recent research has primarily focused on using LLMs to classify text as either human-written and machine-generated texts, our study focus on characterizing these texts using a set of linguistic features across different linguistic levels such as morphology, syntax, and semantics. We select a dataset of human-written and machine-generated texts spanning 8 domains and produced by 11 different LLMs. We calculate different linguistic features such as dependency length and emotionality and we use them for characterizing human-written and machine-generated texts along with different sampling strategies, repetition controls and model release date. Our statistical analysis reveals that human-written texts tend to exhibit simpler syntactic structures and more diverse semantic content. Furthermore, we calculate the variability of our set of features across models and domains. Both human and machine texts show stylistic diversity across domains, with humans displaying greater variation in our features. Finally, we apply style embeddings to further test variability among human-written and machine-generated texts. Notably, newer models output text that is similarly variable, pointing to an homogenization of machine-generated texts.
@@
Sergio E. Zanotto1, Segun Aroyehun2 1Department of Linguistics & Cluster of Excellence “The Politics of Inequality” University of Konstanz sergio.zanotto@uni-konstanz.de 2Department of Politics and Public Administration University of Konstanz segun.aroyehun@uni-konstanz.de
@@ 1 Introduction
The rapid advancements in language models have significantly improved their ability to generate natural language, making machine-generated text (MGT) increasingly indistinguishable from human-written text (HWT). Indeed, recent studies indicate that disinformation produced by state-of-the-art large language models (LLMs) is often perceived as more credible than that created by humans (Spitale et al., 2023). This evolution has highlighted the importance of identifying MGT due to legitimate concerns about the potential for malicious actors to disseminate false information, as well as the broader need to uphold trust and authenticity across online platforms (Chakravarthi et al., 2025; Li et al., 2024; Sarvazyan et al., 2023). Best systems for this task, often referred as Human/Machine Authorship Attribution, all imply the use of LLMs, and different studies show marginal gains of leveraging stylometric features combined with LLMs for successfully achieving this task (Alecakir et al., 2024; Wang et al., 2024a).
In this study, we utilize the RAID dataset (Dugan et al., 2024), a large-scale corpus designed for testing detection tools for machine-generated text and human-written text. This work focuses on analyzing different levels of linguistic analysis, such as morphology, syntax, and semantics, in order to characterize MGT from 11 different models with 4 different decoding strategies, along with the release date of the models. Thus, we analyze a set of representative linguistic features to distinguish human-written text (HWT) from machine-generated text (MGT), potentially uncovering distinctive linguistic patterns in a multi-domain setting (across 8 different domains). Moreover, we focus on linguistic variability over time to check for whether there are patterns of linguistic homogenization in newer LLMs (Sourati et al., 2025; Padmakumar and He, 2023).
We test the following set of linguistic features per level of linguistic analysis: Textual level (Text Length, Sentence Length), Morphology (Morphological Complexity Index for Verbs and Nouns), Syntax (Dependency Tree Depth, Dependency Length), Lexical Level (Word Prevalence, Type-Token Ratio), Semantics (Semantic Similarity), Emotionality. We consider these features as representative for different levels of linguistic analysis of our interest, ensuring overlap with previous work on distinguishing between HWT and MGT (Zanotto and Aroyehun, 2024; Guo et al., 2023; Uchendu et al., 2020). We do not aim to test all possible linguistic features tested in other research on Human/Machine authorship attribution tasks (e.g., Simón et al., 2023; Uchendu et al., 2020).
@@ 2 Related Work
Scholars have explored various approaches to tackle the challenge of distinguishing between human-written and machine-generated texts. This task, often referred as Human/Machine Authorship Attribution (Alecakir et al., 2024), involves detecting whether a text is produced by a human or a generative language model, or attributing authorship among different models.
@@
The Human/Machine attribution of authorship to a text carries significant social relevance, especially in areas such as fake news detection (Kumarage et al., 2023; Jawahar et al., 2020). The need for explainability becomes particularly important when engaging with a broad audience of non-experts, who may not have the means to access or comprehend detection models (Gehrmann et al., 2019). As a result, numerous studies have focused on identifying human-explainable features that can differentiate between machine-generated (MGT) and human-written text (HWT) (e.g., Dugan et al., 2023; Guo et al., 2023; Kumarage et al., 2023; Uchendu et al., 2020). To achieve this, researchers have employed diverse analytical approaches, including stylometric analysis (Kumarage et al., 2023; Ma et al., 2023), qualitative assessments (Guo et al., 2023; Gehrmann et al., 2019), and linguistic feature analysis (Wang et al., 2024b; Uchendu et al., 2020; Ferracane et al., 2017), to diverse corpora, contexts, and generation tasks. Notably, different studies highlight the difficulties encountered by humans in distinguishing machine-generated texts from human-written texts, but little has been done to address the real life consequences of this issue. (e.g., Chakraborty et al., 2023; Jakesch et al., 2023; Fraser et al., 2024).
@@
Moreover, classical machine learning algorithms, such as logistic regression, have been employed to train models on bag-of-words features to differentiate between HWT and MGT (Solaiman et al., 2019; Ippolito et al., 2020). Other traditional methods leverage linguistic features, including POS-tags (Ferracane et al., 2017), topic modeling (Seroussi et al., 2014), and LIWC (Linguistic Inquiry and Word Count) features to provide deeper insights into the characteristics of MGT (Uchendu et al., 2020; Li et al., 2014).
@@
With the advent of Large Language Models (LLMs), fine-tuned models such as RoBERTa have achieved state-of-the-art performance in many tasks (Crothers et al., 2023; Jawahar et al., 2020). Thus, different shared tasks such as SemEval and IberLEF focus on the creation of the best models for tackling Human/Machine Authorship Attribution, with LLMs-based systems always reaching the top positions (Wang et al., 2024a; Sarvazyan et al., 2023). Indeed, classifiers solely based on stylometric features reach poor accuracy in distinguishing Human/Machine authorship (Alecakir et al., 2024; Sharma and Mansuri, 2024), with marginal gains when combined with LLMs (Sharma and Mansuri, 2024). In the case of the Human/machine Authorship Attribution, Transformers encoders leveraging models with token-level probabilistic features offer state-of-the-art detection capabilities (Sarvazyan et al., 2024; Mitchell et al., 2023). Despite providing literature on classifiers for Human/Machine authorship attribution, the focus of our study is to examine possible characteristics that distinguish HWT and MGT on different levels of linguistic analysis. Few studies show that MGT tend to exhibit shorter texts on average, with lower emotional content and syntactically more complex structures compared to HWT (Zanotto and Aroyehun, 2024; Guo et al., 2023). However, they overlook distinctions across models, decoding strategies, or stylistic variability. In doing this, we test variability among models and provide results that possibly point to the tendencies of recent models to exhibit less variability in their output possibly due to the recent practice of training on machine-generated data (Shumailov et al., 2023).
@@ 3 Data
For our research. we use the RAID dataset (Dugan et al., 2024), that is a large-scale corpus comprising over 6.2 million text generations, developed to evaluate detection methods for outputs from language models. The training set of the dataset is constructed from 13,371 human-written documents sampled across eight diverse domains — Abstracts, Books, News, Poetry, Recipes, Reddit, Reviews, and Wikipedia — offering a broad spectrum of human-written and machine-generated English texts for linguistic analysis. For each document, a corresponding generation prompt is dynamically created using either “Chat” or “Non-Chat” templates, ensuring compatibility with the target language model. “Chat” models are fine-tuned (often via supervised fine-tuning and reinforcement learning) on multi-turn interactions with humans. “Non-Chat” models are trained for next-token prediction on unstructured text. Texts are generated using eleven models, namely GPT-2 XL, GPT-3, GPT-4, ChatGPT, Mistral-7B, Mistral-7B (Chat), MPT-30B, MPT-30B (Chat), LLaMA 2 70B (Chat), Cohere, and Cohere (Chat), under four decoding strategies (i.e., greedy decoding and sampling decoding paired with the presence, where possible, or absence of repetition penalties). Greedy models are more deterministic and repetitive, while sampling models introduces randomness to make the response more diverse, whereas the presence of a penalty on repetition avoids excessive looping and redundancy. Note that we exclude texts generated using adversarial attacks from the original dataset. Hence, the selected subset of the dataset is designed to reflect realistic generation scenarios, providing a robust source for analyzing linguistic similarities and differences between machine-generated and human-written texts. In the end, we consider a total amount of 467,985 texts.
@@ 4 Methodology
We extract a range of linguistic features to analyze differences between human-written and machine-generated texts. These features include text length, sentence length, morphological complexity index, dependency length, dependency depth, word prevalence, type-token ratio, semantic similarity, and emotionality. For each feature, we compute the mean and standard error of the mean for human-written and machine-generated texts and assess similarity between human and model distributions using the Mann-Whitney U-test (McKnight and Najab, 2010). To further explore stylistic variability, we apply a style embedding model to extract a representation of the writing style of each text, enabling a comparison of style consistency across human and machine-generated texts. We train a logistic classifier on all extracted linguistic features to predict text authorship based on linguistic features. Additionally, we assess variability across models and domains by computing the standard deviation of the Euclidean distance for all linguistic features across HWT and MGT. We employ Principal Component Analysis (PCA) for dimensionality reduction using style embeddings to quantify variability within models and domains.
@@ 4.1 Features Collection
Text Length: Text Length is calculated by counting the total number of alphanumerical tokens in a text using the tokenizer in SpaCy “en_core_web_sm”. We average the length of all texts produced by humans or models with their own decoding and penalty setting. Text length provides shallow linguistic information on possible differences between humans and models and between models’ generation settings.
Sentence Length: Sentence Length is calculated by counting the total number of alphanumerical tokens per sentence, normalized by the total number of sentences produced by a specific model or human. This measures allows for a better fine-grained understanding of possible differences within text length, whereas two texts could have the same length, but different number of sentences and therefore different sentence length.
@@
Dependency Length: Dependency Length is calculated using the dependency parser in SpaCy “en_core_web_sm”. It represents the distance in number of tokens between a word (dependent) and its syntactic head. We considered for this measure only the intervening words between the head and the last dependent word. This measure focuses on the linear arrangement of words. Longer dependency lengths can indicate that related words are spread farther apart, which reflect more complex syntactic structures.
@@
Style Embeddings: We compute style embeddings using the StyleDistance model (Patel et al., 2025). This model generates style embeddings with a dimension of 768 such that texts with similar stylistic features are closer in embeddings space independent of content. We use these embeddings to assess whether models and humans exhibit different stylistic variability in their generated texts.
@@ 4.2 Models and Domain variability
For each model and domain, we first compute the centroid by taking the mean of the values of our linguistic features. Then, for each individual observation in that model or domain, we calculate the Euclidean distance from its feature vector to the centroid. The overall variability for the models and domains is quantified as the standard deviation of these distances. These standard deviations reflects how spread out the observations are in the feature space relative to the centroid. Moreover, we apply dimensionality reduction using PCA to the style embeddings to match the linguistic feature count (10 dimensions). We compute variability within domains and models for the style embeddings by calculating their centroids, Euclidean distances, and standard deviations. Finally, we map models to their release dates to identify a possible trend in the evolution of models with respect to our set of linguistic features and style embeddings.
@@ 4.3 Feature Importance via a Logistic Classifier
We train a binary logistic regression classifiers with a set of linguistic features to characterize via feature importance analysis human-written texts (label: 1) from machine-generated texts (label: 0). To ensure a balanced dataset, we downsample machine-generated texts, resulting in a training set of 21,376 documents and a test set of 5,344. Rather than comparing with existing methods, our focus is on assessing feature importance to get an overall picture of the impact of these features in characterizing HWT and MGT. We do not aim to build a state-of-the-art detection classifier nor to compare classification results with existing literature.
@@ 5 Results
In this section, we present the linguistic features used to analyze human-written texts (HWT) and machine-generated texts (MGT). These features capture differences in text structure, morphological complexity, syntactic complexity, lexical diversity, and emotionality. We show the results and feature importance from the classifier based on the calculated linguistic features. Section D presents visualizations of our set of linguistics features for each domain. These plots highlight the influence of genre-specific constraints on different domains. In more creative domains such as poetry and books, human-written texts tend to be longer and semantically more diverse. Across all domains, humans consistently use simpler syntactic structures, while maintaining a relatively stable use of common words, as measured by Word Prevalence. Notably, this measure varies among humans between Reddit and news: human texts on Reddit show the highest use of common words, whereas News articles show the lowest.
@@
Table 1: Average values $\pm$ standard errors (SE) of calculated linguistic features per model in the dataset. The pattern of each model’s name is: model + decoding strategy + yes/no penalty on repetitions. Models’ non-bold entries are statistically different to human writing in the respective feature. Bold entries are not statistically different ($p\geq 0.05$), indicating that the feature distribution in the machine-generated text is not distinguishable from human written text. We present graphical versions of this table in the Appendix (Section D)
@@
Table 1 shows the statistical analysis of our selected set of features for HWT and MGT in the entire dataset. Bold entries are those models that are statistically similar to humans for that feature. Different models and decoding strategies manifest different behaviors per feature in comparison to humans. We notice a tendency for chat models and for sampling strategy to output texts that are more similar to humans. Thus, we rely on the feature importance of a logistic classifier to extract further tendencies. Figure 1 illustrates the feature importance for the overall dataset. The classifier primarily relies on simpler syntactic structures (e.g., shorter dependency length and depth) and semantic properties (e.g., lower word prevalence, reduced semantic similarity, and a lower type-token ratio) for distinguishing between human-written and machine-generated texts. We present analysis of each feature in the next sections.
@@
Figure 1: Feature importance for the logistic classifier’s predictions across all domains. A positive coefficient means that as the feature’s value increases, the model is more likely to predict human authorship, while a negative coefficient indicates that higher values make the model less likely to predict human authorship.
5.1 Text and Sentence Length

### v3

@@ Abstract
The rapid advancements in large language models (LLMs) have significantly improved their ability to generate natural language, making texts generated by LLMs increasingly indistinguishable from human-written texts. While recent research has primarily focused on using LLMs to classify text as either human-written or machine-generated texts, our study focuses on characterizing these texts using a set of linguistic features across different linguistic levels such as morphology, syntax, and semantics. We select a dataset of human-written and machine-generated texts spanning 8 domains and produced by 11 different LLMs. We calculate different linguistic features such as dependency length and emotionality, and we use them for characterizing human-written and machine-generated texts along with different sampling strategies, repetition controls, and model release dates. Our statistical analysis reveals that human-written texts tend to exhibit simpler syntactic structures and more diverse semantic content. Furthermore, we calculate the variability of our set of features across models and domains. Both human- and machine-generated texts show stylistic diversity across domains, with human-written texts displaying greater variation in our features. Finally, we apply style embeddings to further test variability among human-written and machine-generated texts. Notably, newer models output text that is similarly variable, pointing to a homogenization of machine-generated texts.
@@
Sergio E. Zanotto1  and Segun Aroyehun2 1Department of Linguistics & Cluster of Excellence “The Politics of Inequality” University of Konstanz sergio.zanotto@uni-konstanz.de 2Department of Politics and Public Administration University of Konstanz segun.aroyehun@uni-konstanz.de
@@ 1 Introduction
The rapid advancements in language models have significantly improved their ability to generate natural language, making machine-generated texts (MGT) increasingly indistinguishable from human-written texts (HWT). Indeed, recent studies indicate that disinformation produced by state-of-the-art large language models (LLMs) is often perceived as more credible than that created by humans (Spitale et al., 2023). This evolution has highlighted the importance of identifying MGT due to legitimate concerns about the potential for malicious actors to disseminate false information, as well as the broader need to uphold trust and authenticity across online platforms (Chakravarthi et al., 2025; Li et al., 2024; Sarvazyan et al., 2023). Best systems for this task, often referred to as Human/Machine Authorship Attribution, all imply the use of LLMs, and different studies show marginal gains of leveraging stylometric features combined with LLMs for successfully achieving this task (Alecakir et al., 2024; Wang et al., 2024a).
In this study, we utilize the RAID dataset (Dugan et al., 2024), a large-scale corpus designed for testing detection tools for machine-generated text and human-written text. This work focuses on analyzing different levels of linguistic analysis, such as morphology, syntax, and semantics, in order to characterize MGT from 11 different models with 4 different decoding strategies, along with the release date of the models. Thus, we analyze a set of representative linguistic features to distinguish human‐written texts (HWT) from machine‐generated texts (MGT), potentially uncovering distinctive linguistic patterns in a multi-domain setting (across 8 different domains). Moreover, we focus on linguistic variability over time to check whether there are patterns of linguistic homogenization in newer LLMs (Sourati et al., 2025; Padmakumar and He, 2023).
We test the following set of linguistic features per level of linguistic analysis: Textual level (Text Length, Sentence Length), Morphology (Morphological Complexity Index for Verbs and Nouns), Syntax (Dependency Tree Depth, Dependency Length), Lexical Level (Word Prevalence, Type-Token Ratio), Semantics (Semantic Similarity), Emotionality. We consider these features as representative of different levels of linguistic analysis of our interest, ensuring overlap with previous work on distinguishing between HWT and MGT (Zanotto and Aroyehun, 2024; Guo et al., 2023; Uchendu et al., 2020). We do not aim to test all possible linguistic features tested in other research on Human/Machine authorship attribution tasks (e.g., Simón et al., 2023; Uchendu et al., 2020).
@@ 2 Related Work
Scholars have explored various approaches to tackle the challenge of distinguishing between human-written and machine-generated texts. This task, often referred to as Human/Machine Authorship Attribution (Alecakir et al., 2024), involves detecting whether a text is produced by a human or a generative language model, or attributing authorship among different models.
@@
The Human/Machine attribution of authorship to a text carries significant social relevance, especially in areas such as fake news detection (Kumarage et al., 2023; Jawahar et al., 2020). The need for explainability becomes particularly important when engaging with a broad audience of non-experts, who may not have the means to access or comprehend detection models (Gehrmann et al., 2019). As a result, numerous studies have focused on identifying human-explainable features that can differentiate between machine-generated (MGT) and human-written texts (HWT) (e.g., Dugan et al., 2023; Guo et al., 2023; Kumarage et al., 2023; Uchendu et al., 2020). To achieve this, researchers have employed diverse analytical approaches, including stylometric analysis (Kumarage et al., 2023; Ma et al., 2023), qualitative assessments (Guo et al., 2023; Gehrmann et al., 2019), and linguistic feature analysis (Wang et al., 2024b; Uchendu et al., 2020; Ferracane et al., 2017), to diverse corpora, contexts, and generation tasks. Notably, different studies highlight the difficulties encountered by humans in distinguishing machine-generated texts from human-written texts, but little has been done to address the real-life consequences of this issue. (e.g., Chakraborty et al., 2023; Jakesch et al., 2023; Fraser et al., 2024).
@@
Moreover, classical machine learning algorithms, such as logistic regression, have been employed to train models on bag-of-words features to differentiate between HWT and MGT (Solaiman et al., 2019; Ippolito et al., 2020). Other traditional methods leverage linguistic features, including POS-tags (Ferracane et al., 2017), surface-level features such as readability indexes or punctuation marks (Doughman et al., 2024; Malviya et al., 2025), topic modeling (Seroussi et al., 2014), sentiment analysis (Hossen Rujeedawa et al., 2025), and LIWC (Linguistic Inquiry and Word Count) features to provide deeper insights into the characteristics of MGT (Uchendu et al., 2020; Li et al., 2014). HWT tend to be longer, express more sentiment polarity, especially negative sentiment, and show greater variability in discourse structures compared to MGT (e.g., Kim et al., 2024; Zanotto and Aroyehun, 2024).
@@
With the advent of Large Language Models (LLMs), fine-tuned models such as RoBERTa have achieved state-of-the-art performance in many tasks (Crothers et al., 2023; Jawahar et al., 2020). Thus, different shared tasks such as SemEval and IberLEF focus on the creation of the best models for tackling Human/Machine Authorship Attribution, with LLM-based systems always reaching the top positions (Wang et al., 2024a; Sarvazyan et al., 2023). Indeed, classifiers solely based on stylometric features reach poor accuracy in distinguishing Human/Machine authorship (Alecakir et al., 2024; Sharma and Mansuri, 2024), with marginal gains when combined with LLMs (Sharma and Mansuri, 2024). Models leveraging Transformers encoders with token-level probabilistic features offer state-of-the-art detection capabilities (Sarvazyan et al., 2024; Mitchell et al., 2023). Despite providing literature on classifiers for Human/Machine authorship attribution, the focus of our study is to examine possible characteristics that distinguish HWT and MGT on different levels of linguistic analysis. Few studies show that MGT tend to exhibit shorter texts on average, with lower emotional content and syntactically more complex structures compared to HWT (Zanotto and Aroyehun, 2024; Guo et al., 2023). However, they overlook distinctions across models, decoding strategies, or stylistic variability. In doing this, we test variability among models and provide results that possibly point to the tendencies of recent models to exhibit less variability in their output, possibly due to the recent practice of training on machine-generated data (Shumailov et al., 2023).
@@ 3 Data
For our research, we use the RAID dataset (Dugan et al., 2024), a large-scale corpus comprising over 6.2 million text generations, developed to evaluate detection methods for outputs from language models. The training set of the dataset is constructed from 13,371 human-written documents sampled across eight diverse domains — Abstracts, Books, News, Poetry, Recipes, Reddit, Reviews, and Wikipedia — offering a broad spectrum of human-written and machine-generated English texts for linguistic analysis. For each document, a corresponding generation prompt is dynamically created using either “Chat” or “Non-Chat” templates, ensuring compatibility with the target language model. All the models use the same specific set of prompts (refer to Section B in Appendix for examples). “Chat” models are fine‐tuned (often via supervised fine‐tuning and reinforcement learning) on multi‐turn interactions with humans. “Non-Chat” models are trained for next‐token prediction on unstructured text. Texts are generated using eleven models, namely GPT-2 XL, GPT-3, GPT-4, ChatGPT, Mistral-7B, Mistral-7B (Chat), MPT-30B, MPT-30B (Chat), LLaMA 2 70B (Chat), Cohere, and Cohere (Chat), under four decoding strategies (i.e., greedy decoding and sampling decoding paired with the presence, where possible, or absence of repetition penalties). Greedy models are more deterministic and repetitive, while sampling models introduce randomness to make the response more diverse, whereas the presence of a penalty on repetition avoids excessive looping and redundancy. Note that we exclude texts generated using adversarial attacks from the original dataset. Hence, the selected subset of the dataset is designed to reflect realistic generation scenarios, providing a robust source for analyzing linguistic similarities and differences between machine-generated and human-written texts. In the end, we consider a total amount of 467,985 texts.
@@ 4 Methodology
We extract a range of linguistic features to analyze differences between human-written and machine-generated texts. These features include text length, sentence length, morphological complexity index, dependency length, dependency depth, word prevalence, type-token ratio, semantic similarity, and emotionality. For each feature, we compute the mean and standard error of the mean for human-written and machine-generated texts and assess similarity between human and model distributions using the Mann-Whitney U-test (McKnight and Najab, 2010). To further explore stylistic variability, we apply a style embedding model to extract a representation of the writing style of each text, enabling a comparison of style consistency across human-written and machine-generated texts. We train a logistic classifier on all extracted linguistic features to predict text authorship based on linguistic features. Additionally, we assess variability across models and domains by computing the standard deviation of the Euclidean distance for all linguistic features across HWT and MGT. We employ Principal Component Analysis (PCA) for dimensionality reduction using style embeddings to quantify variability within models and domains. All codes are available at https://github.com/Sergio-E-Zanotto/LingAIHUman.
@@ 4.1 Features Collection
Text Length: Text Length is calculated by counting the total number of alphanumerical tokens in a text using the tokenizer in SpaCy “en_core_web_sm”. We average the length of all texts produced by humans or models with their own decoding and penalty settings. Text length provides shallow linguistic information on possible differences between humans and models and between different generation settings of models.
Sentence Length: Sentence Length is calculated by counting the total number of alphanumerical tokens per sentence, normalized by the total number of sentences produced by a specific model or human. This measure allows for a better fine-grained understanding of possible differences within text length, whereas two texts could have the same length, but a different number of sentences and therefore different sentence lengths.
@@
Dependency Length: Dependency Length is calculated using the dependency parser in SpaCy “en_core_web_sm”. It represents the distance in number of tokens between a word (dependent) and its syntactic head. We considered for this measure only the intervening words between the head and the last dependent word. This measure focuses on the linear arrangement of words. Longer dependency lengths can indicate that related words are spread farther apart, which reflects more complex syntactic structures.
@@
Style Embeddings: We compute style embeddings using the StyleDistance model (Patel et al., 2025). This model generates style embeddings with a dimension of 768 such that texts with similar stylistic features are closer in embedding space, independent of content. We use these embeddings to assess whether models and humans exhibit different stylistic variability in their generated texts. We selected StyleDistance (Patel et al., 2025) because it is a state-of-the-art model explicitly designed to generate content-independent style embeddings. The model has been shown to generalize well to real-world benchmarks and outperform leading style representation methods across multiple downstream tasks. Evaluations on style identification, style transfer, and authorship verification demonstrate that StyleDistance captures stylistic properties robustly, including under out-of-domain and out-of-distribution conditions. We use the resulting 768-dimensional embeddings as a latent representation of linguistic style, providing an additional analysis alongside our main feature-based approach. While we do not assume a one-to-one correspondence between embedding dimensions and linguistic features, these embeddings allow us to assess whether global stylistic patterns across models and domains align with those revealed through explicit linguistic features.
@@ 4.2 Models and Domain Variability
For each model and domain, we first compute the centroid by taking the mean of the values of our linguistic features. Then, for each individual observation in that model or domain, we calculate the Euclidean distance from its feature vector to the centroid. The overall variability for the models and domains is quantified as the standard deviation of these distances. These standard deviations reflect how spread out the observations are in the feature space relative to the centroid. Moreover, we apply dimensionality reduction using PCA to the style embeddings to match the linguistic feature count (10 dimensions). We compute variability within domains and models for the style embeddings by calculating their centroids, Euclidean distances, and standard deviations. Finally, we map models to their release dates to identify a possible trend in the evolution of models with respect to our set of linguistic features and style embeddings.
@@ 4.3 Feature Importance via a Logistic Classifier
We train a binary logistic regression classifier with a set of linguistic features to characterize via feature importance analysis human-written texts (label: 1) from machine-generated texts (label: 0). To ensure a balanced dataset, we downsample machine-generated texts, resulting in a training set of 21,376 documents and a test set of 5,344. Rather than comparing with existing methods, our focus is on assessing feature importance to get an overall picture of the impact of these features in characterizing HWT and MGT. We do not aim to build a state‑of‑the‑art detection classifier nor to compare classification results with existing literature.
@@ 5 Results
In this section, we present the linguistic features used to analyze human-written texts (HWT) and machine-generated texts (MGT). These features capture differences in text structure, morphological complexity, syntactic complexity, lexical diversity, and emotionality. We show the results and feature importance from the classifier based on the calculated linguistic features. Figure 1 illustrates the variation in linguistic features across domains, showing that MGT exhibit comparatively low variability. Section E in the Appendix presents further visualizations and discussions of observed patterns across domains. These plots highlight the influence of genre-specific constraints on different domains. In more creative domains such as poetry and books, human-written texts tend to be longer and semantically more diverse. Across all domains, human-written texts consistently show simpler syntactic structures, while maintaining a relatively stable use of common words, as measured by Word Prevalence. Notably, this measure varies among HWT between Reddit and news: human-written texts on Reddit show the highest use of common words, whereas News articles show the lowest.
Figure 1: Multi-panel heatmap displaying log-normalized feature standard deviation across different text domains and models
@@
Table 1: Average values $\pm$ standard errors (SE) of calculated linguistic features per model in the dataset. The pattern of each model name is: model + decoding strategy + yes/no penalty on repetitions. Models corresponding to non-bold entries are statistically different from human writing in the respective feature. Bold entries are not statistically different based on the Mann-Whitney U-test ($p\geq 0.05$), indicating that the feature distribution in the machine-generated texts is not distinguishable from human-written texts.
@@
Table 1 shows the statistical analysis of our selected set of features for HWT and MGT in the entire dataset. Bold entries are those models that are statistically similar to humans for that feature. Different models and decoding strategies show different behaviours per feature in comparison to humans. We notice a tendency for chat models and for sampling strategies to output texts that are more similar to humans. Thus, we rely on the feature importance of a logistic classifier to extract further tendencies. Figure 2 illustrates the feature importance for the overall dataset. The classifier primarily relies on simpler syntactic structures (e.g., shorter dependency length and depth) and semantic properties (e.g., lower word prevalence, reduced semantic similarity, and a lower type-token ratio) for distinguishing between human-written and machine-generated texts. We present an analysis of each feature in the next section.
@@
Figure 2: Feature importance for the logistic classifier across all domains. A positive coefficient means that as a feature value increases, the model is more likely to predict human authorship, while a negative coefficient indicates that higher values make the model less likely to predict human authorship.
5.1 How do HWT and MGT differ in terms of linguistic features?